Tony flew his ultralight plane to the next town against a head wind of 10mph; the trip took 2 1/2 hours. The return trip with the same wind took 1 1/2 hours. what was the speed of tonys plane?
Answer by TimothyLamb(4379)

s = d / t
d = s * t
w = 10
---
out: d = (s - w)2.5 = (s - 10)2.5 = 2.5s - 25
bck: d = (s + w)1.5 = (s + 10)1.5 = 1.5s + 15
---
2.5s - 25 = 1.5s + 15
---
s = 40 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php